The decibel rating of a sound is given by: D(i)=10 times log{i/i
0}

where L is intensity of the sound and i0 is the minimum intensity detectable by the human ear. Find the decibel rating of eac of the following sounds:
Rock music 895,000,000,000 i0
Average sound level in the movie king kong 3,600,000,000 i0

Answer by stanbon(48558)

D(L)=10 times log{L/i0}

where L is intensity of the sound and i0 is the minimum intensity detectable by the human ear. Find the decibel rating of each of the following sounds:
Rock music 895,000,000,000 i0
D(8.95x10^11) =10*log (8.95x10^11*i(o)/i(o)]
10*11.951823...
=119.51923.. dB
-----------------
Average sound level in the movie king kong 3,600,000,000 i0
D(3.6X10^9) = 10 log (3.6x10^9*i(o)/i(o)]
= 10*9.5563025...
= 95.563025.. dB
=============
Cheers,
Stan H.